An express train and a charter bus leave Chicago to travel to Champaign. The express train can make the trip in two hours and the bus takes five hours for the trip. The speed of the express train is 42 miles per hour faster than the speed of the bus. Find the speed of the bus?

Answer by ikleyn(52908)

.

Let x be the speed of the bus, in miles per hour.

Then the speed of the train is (x+42) miles per hour.


Then you write this "distance" equation

    5x = 2*(x + 42)


saying that the distance is the same.


Simplify and solve

    5x = 2x + 84

    5x - 2x = 84

    3x      = 84

     x      = 84/3 = 28.


ANSWER.  The bus speed is 28 miles per hour.

Solved.